OpenStudy (mhchen):

X~bin(4,0.6) Probability of x = 0

OpenStudy (mhchen):

I used the \[P(X = x) = \left(\begin{matrix}n \\ x\end{matrix}\right)p^{x}(1-p)^{n-x}\] Where p = probability of success n = number of trials

OpenStudy (mhchen):

So according to this, if x = 0, I should get 0.01024. But the answer is 0.0256

OpenStudy (mhchen):

I hate myself